The length of a rectangle solid is 10. The width, w and the height, h follow the relationship, and w^3h^3= 27. What is the volume of the rectangular solid?
A) 279
B) 30
C) 90
D)60

Answer by greenestamps(13200)



V = lwh

w^3h^3 = (wh)^3 = 27. Solve that to find wh. Then multiply by the given value of l.